OpenStudy (anonymous):

A can of chicken stock measures 3.5 inches in diameter and is 5 inches tall. What is the volume of the can? Use 3.14 for pi. Round your answer to the nearest hundredth. 38.50 cubic inches 48.08 cubic inches 92.46 cubic inches 192.33 cubic inches

OpenStudy (anonymous):

\(V = \pi r^2 h\) \(V = (3.14)(1.75^2)(5)\) Diameter is 3.5, so the radius is half which is 1.75. First multiply 1.75 * 1.75

OpenStudy (anonymous):

ok so it = 3.0625

OpenStudy (anonymous):

Yes, so we have: \(V = (3.14)(3.0625)(5)\) So multiply 3.14 * 3.0625 * 5

OpenStudy (anonymous):

i got 48.08125

OpenStudy (anonymous):

Yep, that's our answer.

OpenStudy (anonymous):

Thanks for the help again!